OpenStudy (anonymous):

evaluate ∫_(-2)^3〖m^3 〖 (5+ m^4)〗^7 dm〗 evaluate ∫_(-2)^3〖m^3 〖 (5+ m^4)〗^7 dm〗 @Mathematics

OpenStudy (anonymous):

\[\int\limits_{-2}^{3}〖m^3 〖 (5+ m^4)〗^7 dm〗\]

OpenStudy (anonymous):

let u=5+m^4 -> du= 4m^3dx x=-2 -> u1= ... x=3 -> u2=... 1/4 limit from u1 to u2 integration of u^7du

OpenStudy (anonymous):

what about the m^3

OpenStudy (anonymous):

do u see m^4 is on ur problem, that is du

OpenStudy (anonymous):

i meant m^3

OpenStudy (anonymous):

du=4m^3dx

OpenStudy (anonymous):

are you familiar with wolfram alpha

OpenStudy (anonymous):

it shows something completely diff

OpenStudy (anonymous):

no

OpenStudy (anonymous):

okay, i think u get lost or don't really understand the basic of deri or inter. I'll show it clearly, let me get my calculator.. l0l

OpenStudy (anonymous):

I'm eating

OpenStudy (anonymous):

Okay, let u=5+m^4 -> du= 4m^3dx m=-2 -> u= 5+ (-2)^4= 21 m=3 -> u= 5 +(3)^4=86 \[1/4\int\limits_{21}^{86} u^7du\] =(1/4)u^8/8 from 21 to 86 Then we have: [(86)^8] / 32 - [(21)^8] / 32 = ...

OpenStudy (anonymous):

Find the answer and replace to find m, that's all

OpenStudy (anonymous):

how did you get 56 21definite integral is all i want to know

OpenStudy (anonymous):

thanks for the rest

OpenStudy (anonymous):

i don't have 56 on my answer

OpenStudy (anonymous):

look at m and u, thats how i found the limit

OpenStudy (anonymous):

\[\int\limits _{-2}^3m^3 \left(5+m^4\right)^7dm=\frac{2992141448206495}{32}=93504420256452 \]